Two cities, A and B, are connected by means of a highway 150 kms long. An automobile leaves A at
1:00 pm and travels at a uniform rate of 40 kms per hour toward B. Thirty minutes later, another
automobile leaves city A and travels toward B at a uniform rate 55 kms per hour. At what time will
the second car overtake the first car?

Answer by ikleyn(52792)

.
Two cities, A and B, are connected by means of a highway 150 kms long. An automobile leaves A at
1:00 pm and travels at a uniform rate of 40 kms per hour toward B. Thirty minutes later, another
automobile leaves city A and travels toward B at a uniform rate 55 kms per hour. At what time will
the second car overtake the first car?
~~~~~~~~~~~~~~~~~~~

Let t be the time after 1:00 pm, in hours.


The distance traveled by first auto is 40t kilometers.

The distance traveled by the second auto is 55*(t-0.5) kilometers.


We want to find t when the distance is the same (overtake condition)

    40t = 55(t-0.5).


Simplify this equation and find t

    40t = 55t - 0.5*55

     0.5*55 = 55t - 40t

     27.5 = 15t

     t = 27.5/15 = 55/30 = 110/60 = 1 hour and 50 minutes.


The second auto will overtake the first at 1 hour and 50 minutes after 1:00 pm, which is 2:50 pm.


The traveled distance is less than 40 km/h * 2 hours = 80 kilometers;
hence, the second car overtakes the first one before than either car will get the city B.

Solved.

-----------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.